OpenStudy (johnt):

I'm trying to understand limits at infinity better. This problem is giving me fits: limit at negative infinity of (x + sqrt(x^2 + 2x + 3))

OpenStudy (johnt):

Here is a propperly drawn equation: \[\lim_{x \rightarrow -\infty} (x + \sqrt{x^{2} + 2x + 3})\]

OpenStudy (anonymous):

use componendo dividendo

OpenStudy (johnt):

I do not understand what you mean.

OpenStudy (anonymous):

gimmick is to multiply by \[\frac{x-\sqrt{x^2+2x+3}}{x-\sqrt{x^2+2x+3}}\]

hartnn (hartnn):

or u can do 2 substitutions: first, put x by -x and make x-> infinity then put x = 1/y to make y->0

OpenStudy (anonymous):

i mean to say if you do it like this \[x+\sqrt{x ^{2}+2x+3}=\frac{x+ \sqrt{(x ^{2}+2x+3) }+1 }{x+\sqrt{x ^{2}+2x+3}-1 }\] take x common and then apply the limits

OpenStudy (johnt):

"x common" ??

OpenStudy (johnt):

Oh, and BTW at my current point in the course, we're not supposed to use L'Hopital's rule. There is a possibility my TA made a mistake putting this in the study guide if this requires L'Hopital's rule to solve.

OpenStudy (zarkon):

it does not require L'Hospitals rule

OpenStudy (zarkon):

I would start with satellite73's 'gimmick'

OpenStudy (johnt):

I did. And I get \[\lim_{x \rightarrow -\infty} \frac{ -2x - 3 }{ x - \sqrt{x^2 + 2x + 3} }\]

OpenStudy (johnt):

My problem comes because I do not understand what to do with \[\sqrt{-\infty^2 + 2(-\infty) + 3}\]

OpenStudy (zarkon):

factor out an \(x^2\) from under the radical

OpenStudy (anonymous):

by my suggested procedure you need to divide both numerator and denominator with x

OpenStudy (johnt):

@Zarkon would that then basically cause \[\sqrt{2(- \infty) + 3}\] to be ignorable?

OpenStudy (anonymous):

my answer is 2

OpenStudy (zarkon):

the answer is not 2

OpenStudy (johnt):

@harsh314 the correct answer is \[-\infty \] I just don't know how to get there.

OpenStudy (zarkon):

are you sure it is \(-\infty\)

OpenStudy (zarkon):

I get -1

OpenStudy (johnt):

Yes. That is what my TA provided as the answer.

OpenStudy (anonymous):

sorry i did a small mistake

OpenStudy (zarkon):

that is not correct

OpenStudy (zarkon):

it is -1

OpenStudy (zarkon):

wolfram agrees with me... http://www.wolframalpha.com/input/?i=Limit [x%2Bsqrt%28x^2%2B2x%2B3%29%2Cx%2C-infinity]

OpenStudy (johnt):

Oh! You're right. http://www.wolframalpha.com/input/?i=limit+at+-+infinity+%28x+%2B+sqrt%28x%5E2+%2B+2x+%2B+3%29%29 Well crap, now what do I do?

OpenStudy (zarkon):

type Limit[x+sqrt(x^2+2x+3),x,-infinity] into wolfram alpha

OpenStudy (zarkon):

do what I said above...factor out an \(x^2\) from under the radical

OpenStudy (anonymous):

but i got two answers one is 3 and the other is minus one due to the square root

OpenStudy (zarkon):

using the usual topology you can only get one answer for a limit

OpenStudy (johnt):

Ah, @Zarkon I do get -1 now. You never answered my question above though-- the reason that works is because it causes the square root in the denominator to be ignorable?

OpenStudy (anonymous):

it is not exacty compulsory to get one answer only plz try it by the one i suggested

OpenStudy (johnt):

\[\frac{ -2(- \infty) - 3 }{ -\infty - -\infty \sqrt{2(-\infty) - 3} }\]

OpenStudy (johnt):

That would simplify to:\[\frac{ 2\infty }{ -2\infty }\]

OpenStudy (anonymous):

the reason why iam saying so is if you try to put a limit of x tending to 1 on the function\[\sqrt{x}\] then you get two answers one is -1 and the other is+1

OpenStudy (zarkon):

\[\lim_{x \to-\infty} \frac{ -2x - 3 }{ x - \sqrt{x^2 + 2x + 3} }\] \[=\lim_{x \to-\infty} \frac{ -2x - 3 }{ x - \sqrt{x^2}\sqrt{x + \frac{2}{x} + \frac{3}{x^2}} }\] \[=\lim_{x \to-\infty} \frac{ -2x - 3 }{ x - |x|\sqrt{x + \frac{2}{x} + \frac{3}{x^2}} }\] our x<0 (it is going to -infinity so |x|=-x \[=\lim_{x \to-\infty} \frac{ -2x - 3 }{ x +x\sqrt{x + \frac{2}{x} + \frac{3}{x^2}} }\] factor out the x in the bottom ...simplify...take limit

OpenStudy (zarkon):

small tyop

OpenStudy (zarkon):

typo ;)

OpenStudy (johnt):

My knowledge breaks down at \[\sqrt{-\infty}\] Could you please explain what happens to that?

OpenStudy (zarkon):

\[=\lim_{x \to-\infty} \frac{ -2x - 3 }{ x+x\sqrt{1 + \frac{2}{x} + \frac{3}{x^2}} }\]

OpenStudy (johnt):

Ok, @Zarkon I see your correction. So, you would have the "square root of 1 - very small - very small"? Which would then be 1?

OpenStudy (zarkon):

yes..\(\sqrt{1}=1\)

OpenStudy (zarkon):

\[=\lim_{x \to-\infty} \frac{ -2x - 3 }{ x(1+\sqrt{1 + \frac{2}{x} + \frac{3}{x^2}} )}\] multiply top and bottom by 1/x \[=\lim_{x \to-\infty} \frac{ -2 - (3/x) }{ 1+\sqrt{1 + \frac{2}{x} + \frac{3}{x^2}} }\] \[=\frac{-2+0}{1+1}=-1\]

OpenStudy (johnt):

OK, I think I understand now. Thank you so much!

OpenStudy (zarkon):

no problem